Clearly explain it to me.
Factor f(x) into linear factors given that k is a zero of f(x). f(x) = 3x^3 + x^2 - 82x + 120; k = 4
Answer #1

foy 3x3+x2-5277120 Crinen K= 4 es of f(21) (-4) is a linear factor of the a zirro zuroby synth hetic devision 4 3 82 52 120 -120 12 13 w o-137-3) Now Then flay (1-4) Зи-- Зу-t13-3 3-52-8х - 30 = 7 (3х-5)-( 3х-5) (37-5)(-6) fon) ----u(-6) 15м- S г. г/ и

Hope it helps.